OpenStudy (theedwardsfamily):

Solve. 8 – 3x – x = 12 x =

OpenStudy (anonymous):

\[8-4x=12\]

OpenStudy (anonymous):

subtract 8 on both sides

OpenStudy (anonymous):

\[8-8-4x=12-8\]

OpenStudy (anonymous):

\[-4x=4\]

OpenStudy (theedwardsfamily):

ok

OpenStudy (anonymous):

divide -4 on both sides

OpenStudy (anonymous):

\[x=?\]

OpenStudy (anonymous):

solve it @TheEdwardsFamily

OpenStudy (theedwardsfamily):

@MARC_ my answer is -1

OpenStudy (anonymous):

correct :)

OpenStudy (theedwardsfamily):

thnx

OpenStudy (anonymous):

yw